DISPERSED FLUORSCENCE SPECTROSCOPY OF SOME ALKOXY RADICALS IN A SUPERSONIC JET

\underline{JIN JIN}, ANDREW BEZANT and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 West 18th Avenue, Columbus, Ohio 43210.

The dispersed fluorescence spectra of the B̃ − X̃ transition of some alkoxy radicals have been obtained in a supersonic jet environment. The alkoxy radicals were produced by tripled Nd: YAG (355 nm) photolysis of the corresponding alkyl nitrite structural isomers. Fluorescence in the near ultraviolet from levels of the B̃ state of the alkoxy radical, in a specific conformation, has been dispersed by a monochromator with a resolution of 0.1nm and recorded by an ICCD camera. Analysis of the spectra yield information about the vibronic structure of the ground state of the radical.